What is the chance of getting at least one defective item if 3 items are drawn randomly from a lot containing 6 items of which 2 are defective items?
Answer by stanbon(75887)

What is the chance of getting at least one defective item if 3 items are drawn randomly from a lot containing 6 items of which 2 are defective items?
-------------------
n = 3 ; number of defective = 2 ; number of not defective = 4
-----
P(at least one defective) = 1 - P(3 not defective)
----
= 1 - 4C36C3
----
= 1 - 4/20
---------
= 1 - 1/5
------
= 4/5
--------------------
Cheers,
Stan H.
=========================